# Equilibria in non-Euclidean geometries

### v1

@@ 4.2. The case $\mathbb{V}^{3}=\mathbb{H}^{3}$
For any point $x\in\mathbb{R}^{3}$, let $r(x)$ denote the Euclidean distance of $x$ from the origin $o$. Then the following holds.
@@
    $d_{H}(o,x)=\operatorname{arctanh}r(x)=\frac{1}{2}\ln\frac{1+r(x)}{1-r(x)}.$
@@
    $\mu_{3}(X)=\int_{x\in g_{N}(X)}\frac{1}{\left(1-\left(r(x)\right)^{2}\right)^{2}}\,d\lambda_{3}(x)$
@@
(4.10) $M_{H}(K)=\int_{x\in K}\frac{x_{3}}{\left(1-\left(r(x)\right)^{2}\right)^{5/2}}d\lambda_{3}(x).$
@@
Consider any point $x=(x_{1},x_{2},x_{3})\in\mathbb{H}^{3}$, and let $y=(x_{1},x_{2},0)$. By (i) of Lemma 4.9, $\cosh d_{H}(o,x)=\frac{1}{\sqrt{1-(r(x))^{2}}}$ and $\cosh d_{H}(o,y)=\frac{1}{\sqrt{1-(r(y))^{2}}}$. By the hyperbolic Pythagorean Theorem, $\cosh d_{H}(y,x)=\frac{\cosh d_{H}(o,x)}{\cosh d_{H}(o,y)}=\frac{\sqrt{1-(r(y))^{2}}}{\sqrt{1-(r(x))^{2}}}$. Thus, $\sinh d_{H}(y,x)=\sqrt{\cosh^{2}d_{H}(o,x)-1}=\frac{x_{3}}{\sqrt{1-(r(x))^{2}}}$. Combining this formula with the volume element appearing in (ii) of Lemma 4.9, the assertion follows. ∎
Consider a ball $B$ of radius $0<R<1$ centered at $o$. This ball represents a hyperbolic ball of radius $\operatorname{arctanh}R$. In our construction we find a slight deformation $K$ of $B$ such that its (hyperbolic) centroid is $o$, and it has only two equilibrium points. The construction, using the same family $\mathcal{F}_{m}$ defined in Subsection 4.1 in $H_{N}$, is entirely analogous to the one in Subsection 4.1, setting $S=\mathbb{S}^{3}$, and thus, we omit it.
@@ 4.3. The case when $\mathbb{V}^{3}$ is a rotationally symmetric $3$-dimensional normed space
Let $M$ be the unit ball of $\mathbb{V}^{3}$. Without loss of generality, we may assume that $M$ is symmetric to the $x_{3}$-axis of the space. We make use of the ambient Euclidean structure of the space, and denote by $\rho_{M}(u):\mathbb{S}^{2}\to\mathbb{R}$ the radial function of $M$, defined as
@@
For any $0<c>1$ and $0<d<1$, the only equilibrium points of $K(c,d)$ with respect to $o$ are $R_{c,d}(u_{N})u_{N}$ and $R_{c,d}(u_{S})u_{S}$, where $u_{N}=(0,0,1)\in\mathbb{S}^{2}$ and $u_{S}=(0,0,-1)\in\mathbb{S}^{2}$.
@@
Note that by the symmetry of $K(c,d)$, the points $R_{c,d}(u_{N})u_{N}$ and $R_{c,d}(u_{S})u_{S}$ are equilibrium points of $K(c,d)$. Assume that at the parameter value $(\theta_{0},\varphi_{0})$, with $\frac{\pi}{2}<\theta_{0}<\frac{\pi}{2}$, $K(c,d)$ has an equilibrium point, implying that the tangent plane of $K(c,d)$ at this point is Birkhoff orthogonal to $u(\theta_{0},\varphi_{0})$. Let $(\theta,\varphi)\mapsto\overline{R}(\theta,\varphi)$ denote this plane, parametrized with the same parameters $\theta,\varphi$ in a neigborhood of $(\theta_{0},\varphi_{0})$. Then, from
$\left(R_{c,d}\right)^{\prime}_{\theta}(\theta_{0},\varphi_{0})=R\left(\rho_{M}\right)^{\prime}_{\theta}(\theta_{0},\varphi_{0})(1+d\rho_{c}(\theta_{0},\varphi_{0}))+R\rho_{M}(\theta_{0},\varphi_{0})d(\rho_{c})^{\prime}_{\theta}(\theta_{0},\varphi_{0})$
we obtain that $(\rho_{c})^{\prime}_{\theta}(\theta_{0},\varphi_{0})=0$. Similarly, by comparing the partial derivatives of the same two functions with respect to $\varphi$, it follows that $(\rho_{c})^{\prime}_{\varphi}(\theta_{0},\varphi_{0})=0$. Hence, by [11, Lemma 4], we have a contradiction. ∎

### v2

@@ 4.2. The case $\mathbb{V}^{3}=\mathbb{H}^{3}$
For any point $x\in\mathbb{R}^{3}$, let $|x|$ denote the Euclidean distance of $x$ from the origin $o$. Then the following holds.
@@
    $d_{H}(o,x)=\operatorname{arctanh}|x|=\frac{1}{2}\ln\frac{1+|x|}{1-|x|}.$
@@
    $\mu_{3}(X)=\int_{x\in g_{N}(X)}\frac{1}{\left(1-|x|^{2}\right)^{2}}\,d\lambda_{3}(x)$
Lemma 4.10 is the hyperbolic counterpart of Lemma 4.2.
@@
(4.10) $M_{H}(K)=\int_{x\in K}\frac{x_{3}}{\left(1-|x|^{2}\right)^{5/2}}d\lambda_{3}(x).$
@@
Consider any point $x=(x_{1},x_{2},x_{3})\in\mathbb{H}^{3}$, and let $y=(x_{1},x_{2},0)$. By (i) of Lemma 4.9, $\cosh d_{H}(o,x)=\frac{1}{\sqrt{1-|x|^{2}}}$ and $\cosh d_{H}(o,y)=\frac{1}{\sqrt{1-|y|^{2}}}$. By the hyperbolic Pythagorean Theorem, $\cosh d_{H}(y,x)=\frac{\cosh d_{H}(o,x)}{\cosh d_{H}(o,y)}=\frac{\sqrt{1-|y|^{2}}}{\sqrt{1-|x|^{2}}}$. Thus, $\sinh d_{H}(y,x)=\sqrt{\cosh^{2}d_{H}(o,x)-1}=\frac{x_{3}}{\sqrt{1-|x|^{2}}}$. Combining this formula with the volume element appearing in (ii) of Lemma 4.9, the assertion follows. ∎
Consider a ball $B$ of radius $0<R<1$ centered at $o$. This ball represents a hyperbolic ball of radius $\operatorname{arctanh}R$. In our construction we find a slight deformation $K$ of $B$ such that its (hyperbolic) centroid is $o$, and it has only two equilibrium points. The construction, using the same family $\mathcal{F}$ defined in Subsection 4.1 in $H_{N}$, is entirely analogous to the one in Subsection 4.1, setting $S=\mathbb{S}^{3}$, and thus, we omit it.
@@ 4.3. The case when $\mathbb{V}^{3}$ is a rotationally symmetric $3$-dimensional normed space
Let $M$ be the unit ball of $\mathbb{V}^{3}$. Without loss of generality, we may assume that $M$ is symmetric to the $x_{3}$-axis of the space. We make use of the ambient Euclidean structure of the space, and denote by $\rho_{M}(u):\mathbb{S}^{2}\to\mathbb{R}$ the radial function of $M$ [37], defined as
@@
For any $0<c<1$ and $0<d<1$, the only equilibrium points of $K(c,d)$ with respect to $o$ are $R_{c,d}(u_{N})u_{N}$ and $R_{c,d}(u_{S})u_{S}$, where $u_{N}=(0,0,1)\in\mathbb{S}^{2}$ and $u_{S}=(0,0,-1)\in\mathbb{S}^{2}$.
@@
Note that by the symmetry of $K(c,d)$, the points $R_{c,d}(u_{N})u_{N}$ and $R_{c,d}(u_{S})u_{S}$ are equilibrium points of $K(c,d)$. Assume that for some $\frac{\pi}{2}<\theta_{0}<\frac{\pi}{2}$ and setting $u_{0}=(\cos\theta_{0}\cos\varphi_{0},\cos\theta_{0}\sin\varphi_{0},\sin\theta_{0})\in\mathbb{S}^{2}$, $K(c,d)$ has an equilibrium point at $R_{c,d}(u_{0})u_{0}\in\operatorname{bd}(K(c,d))$. Then the tangent plane $H$ of $K(c,d)$ at this point is Birkhoff orthogonal to $u_{0}$. In other words, $H$ is tangent at $R_{c,d}(u_{0})u_{0}$ not only to $K(c,d)$, but also to $R(1+d\rho_{c}(u_{0})M$, i.e. the homothetic copy of $M$ centered at $o$ and containing $R_{c,d}(u_{0})u_{0}$ in its boundary. Thus, we have $\left(R_{c,d}\right)^{\prime}_{\theta}(u_{0})=R(1+d\rho_{c}(u_{0})\left(\rho_{M}\right)^{\prime}_{\theta}(u_{0})$, and $\left(R_{c,d}\right)^{\prime}_{\varphi}(u_{0})=R(1+d\rho_{c}(u_{0})\left(\rho_{M}\right)^{\prime}_{\varphi}(u_{0})$. By the definition of $R_{c,d}$, from this we obtain that $(\rho_{c})^{\prime}_{\theta}(u_{0})=(\rho_{c})^{\prime}_{\varphi}(u_{0})=0$. Hence, by the Euclidean result in [11, Lemma 4], we have a contradiction. ∎